OpenStudy (anonymous):

Write y = –2/3x + 7 in standard form using integers. 2x + 3y = 21 3x – 2y = 21 –2x – 3y = 21 –2x + 3y = 7

OpenStudy (anonymous):

The standard form looks like: Ax + By = C

OpenStudy (anonymous):

So, a good first step would be to get your initial equation in a form that doesn't have any fractions... y = –2/3x + 7 Try multiplying everything by 3.... what is your new equation then?

OpenStudy (anonymous):

Is the answer to the problem A?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

When I solve it, I get: y = –2/3x + 7 3y = -2x + 21 (Multiply through by 3) 2x + 31 = 21 (move 2x to the left side) That matches answer A...

OpenStudy (anonymous):

oops, typo above.... last line should have been 2x + 3y = 21 not sure why it typed "31" where I meant "3y".